OpenStudy (anonymous):

An upward force of 600 N acts on a 50 kg dumbwaiter. The dumbwaiter's acceleration is?

OpenStudy (anonymous):

Well, the net force is 600 N - F(gravity) = 600 - 10*50 = 100 N (letting g approximately equal 10 m/s^2). Therefore, the acceleration is \[a = \frac{F_{net}}{m} = \frac{100 N}{50 kg} = 2 \frac{m}{s^2} \]